Ticket Number: HEA-NOC/20071022-4 Ticket Status: CLOSE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20071022 10:51 UTC+1 Problem Start: 20071022 10:00 UTC+1
Ticket Closed: 20080207 12:16 UTC Problem End: 20071031 13:20 UTC
Site/line: GEANT, JANET, ADVA-NB, Citywest PoP
A new DC rectifier is being installed in our Citywest PoP to replace our
old unsupported kit. The installation is taking place this morning (Mon
22nd Oct) and supplies are due to be switched over on Wednesday morning.
Brief loss of power (~20 each) to the following DC powered kit:
- HEAnet ADVA optical equipment (*2), lighting the NB dark fibre network.
- GEANT optical equipment.
- JANET optical equipment.
20071022 10:40 GM Installation of the rectifier is currently taking place.
20071022 15:25 GM Rectifier installed and GEANT router powered up. Other
devices will be migrated on the morning of Wednesday 31st Oct.
20071026 10:28 GM A maintenance window has been scheduled from 0900 until
1300 on Wednesday 31st October to allow for the migration of DC power
suplies to the new rectifier. Redundant power feeds will be migrated
seperately which should ensure no power outage to any devices.
20071030 15:35 GM The DC migration works are going ahead tomorrow during
the scheduled maintenance window.
20071031 16:20 GM All DC power within our CWT PoP has now been migrated
across to our new rectifier system.
20071122 09:24 GM There are issues with the SNMP monitoring module
currently installed in the rectifier unit, this is being investigated by
20080207 12:02 GAR New SNMP module installed and running, as per ticket
This ticket can be monitored at http://www.hea.net/tickets/20071022-4
HEAnet Limited. Registered in Ireland, No. 275301.